# Controlling Large Language Model-based Agents for Large-Scale Decision-Making: An Actor-Critic Approach

### 09 Dec 2023 v2

@@ 4.1.1 Experimental Settings
System resource allocation HolmesParker et al. (2014) can be viewed as a single-step decision and optimization problem that require mathematical reasoning capabilities of LLMs. It has numerous practical applications, such as addressing traffic congestion. In this context, the primary objective is to achieve effective system resource allocation among multiple traffic controllers acting as agents. These agents play a crucial role in directing vehicles onto the main road. The primary objective entails the collaboration of agents to optimize the utilization of the main route while mitigating congestion, and the ultimate goal is to formulate an optimal resource allocation strategy.
@@
In this scenario, each agent is capable of selecting an integer between 0 and 9 as their action, with no knowledge of the choices made by other agents. Centralized critic possesses the authority to access the actions taken by all agents and the corresponding average values of these actions. The objective for the agents is to synthesize their experiences from multiple decision rounds and infer the allocation scheme that leads to the maximum rewards. This problem can be viewed as an optimization problem, and it does not require additional veracity scrutiny or explicit plan confirmation processes. This particular scenario is highly suited for validating the capabilities of the TripletCritic.
@@ 4.1.2 Results
Table 2: Evaluation results under different grid settings in the Grid Transportation-Easy scene.
[…]
HMAS-2 100% 9.9(2.74) 3.3(2.05) 49.9(17.98)
2x2 LLaMAC 100% 7.0(1.79) 2.0(1.26) 23.9(8.38)
HMAS-2 80% 15.5(6.09) 12.3(5.83) 158.4(107.84)
2x4 LLaMAC 100% 7.6(1.36) 4.3(1.42) 38.0(10.57)
HMAS-2 60% 30.6(9.70) 26.1(13.59) 599.3(245.40)
4x8 LLaMAC 100% 12.9(2.70) 10.7(3.35) 122.6(30.55)
Table 3: Evaluation results under different grid settings in the Grid Transportation-Hard scene.
[…]
HMAS-2 80% 7.0(5.0) 6.0(9.74) 76.1(116.66)
2x2 LLaMAC 100% 4.7(1.35) 3.6(2.80) 28.8(18.49)
HMAS-2 20% 17.0(9.0) 24.0(20.0) 355.5(291.05)
2x4 LLaMAC 90% 7.44(2.95) 10.56(7.54) 94.0(68.09)
HMAS-2 0% - - -
4x8 LLaMAC 90% 8.44(1.57) 12.11(2.51) 119.8(32.75)
As shown in Figure 5, it is evident that within a limited number of steps, LLaMAC demonstrates the ability to explore and learn through continuous interaction with the environment. The final performance of all methods is presented in Figure 6. The TripletCritic approach within LLaMAC exhibits a similar structure to the Multi-agent Debate method, and compared to other approaches, these two methods display relatively stable performance. However, debate-based methods often suffer from excessive or insufficient exploration, resulting in a tendency to converge to local optima. On the other hand, approaches that emphasize exploration and exploitation struggle to maintain stable performance. The former exhibits significant oscillations due to excessive exploration, while the latter prematurely converges to local optima after only a few simple exploratory steps, aligning with the expected characteristics of these methods. Distributed approaches incur the highest access cost, and we limited our testing to scenarios with a maximum of 20 agents, where each agent selects actions based on local observations, failing to capture true relationships.
@@ 4.2 Grid Transportation
In this environment, the Veracity Scrutiny within the Internal Feedback involves policy checks of the joint strategy and is set to evaluate (1) whether the output grammar conforms to the specified format and (2) whether the joint actions result in conflicts. The Plan Confirmation within the External Feedback involves policy checks specific to each agent and is set to evaluate (1) the availability of actions and (2) whether the suggestions result in a shorter Manhattan distance between objects and targets.
We conduct a comparative analysis between our method and the state-of-the-art solution, HMAS-2 Chen et al. (2023b). For each scenario, we conduct tests on grid configurations of $2\times 2$, $2\times 4$, and $4\times 8$, respectively. Table 2 and Table 3 present a comprehensive performance comparison between the two methods, clearly demonstrating the overall superiority of our approach. In complex scenarios involving long-term iterative decision-making, LLaMAC exhibits a significantly higher success rate compared to HMAS-2. Furthermore, LLaMAC consistently achieves task completion in fewer interaction steps, highlighting the performance advantages of its employed strategies. Additionally, the TripletCritic facilitates the generation of superior initial suggestions, thereby reducing the need for feedback iterations and greatly enhancing token utilization efficiency.
Furthermore, during the experimental process, we observe that LLaMAC effectively enhances the capabilities of LLM in long-term planning and execution, spatial reasoning, and learning from interactions or errors. For example, spatial reasoning poses a significant challenge for LLMs, as they are more prone to hallucinations when determining whether an object is closer to the target. This issue becomes more pronounced in the Hard scenario. As shown in Figure 8, in the HMAS-2 method, agents often move objects to positions far from the target and may repeatedly move them between two particular locations. In contrast, in LLaMAC, such occurrences are often corrected during the external feedback phase. The actor only needs to focus on its own task, and when it receives suggestions from the critic, the difficulty of determining the effectiveness of individual agent tasks is significantly reduced compared to joint policies. This makes spatial reasoning errors more easily detected, reflected and corrected.
@@ 5 Conclusion
In this study, we present a novel framework called LLaMAC to enhance the collaborative performance of large-scale multi-agent systems based on Large Language Models. Building upon the commonsense reasoning capabilities exhibited by LLMs, we effectively augment the planning and coordination abilities among agents through stable reasoning mechanisms and comprehensive feedback mechanisms, facilitating continuous interaction between agents and the environment. LLaMAC demonstrates remarkable performance in coordinated scenarios involving a large number of agents. Notably, it exhibits exceptional capabilities in long-term planning, mathematical reasoning and optimization problems, spatial reasoning, and learning from mistakes. Additionally, LLMs reduces the access costs associated with large-scale multi-agent collaboration. We believe that with further enhancements in LLMs and the emergence of more collaboration frameworks, the field of multi-agent collaboration will experience new opportunities for advancement.

### 23 Jan 2024 v3

@@ 4.1.1 Experimental Settings
System resource allocation HolmesParker et al. (2014) can be viewed as a single-step decision and optimization problem that require mathematical reasoning capabilities of LLMs. It has numerous practical applications, such as addressing traffic congestion. In this context, the primary objective is to achieve effective system resource allocation among multiple traffic controllers acting as agents. These agents play a crucial role in directing vehicles onto the main road, optimizing the utilization of the main route while mitigating congestion.
@@
In this scenario, each agent is capable of selecting an integer between 0 and 9 as their action, with no knowledge of the choices made by other agents. The objective for the agents is to synthesize their experiences from multiple decision rounds and infer the allocation scheme that leads to the maximum rewards. Centralized critic possesses the authority to access the actions taken by all agents and the corresponding average values of these actions. This particular scenario is highly suited for validating the capabilities of the TripletCritic.
@@ 4.1.2 Results
Table 2: Evaluation results under different grid settings in the Grid Transportation scenarios include metrics such as the success rate (Success), time steps (Steps) taken to execute tasks, and the count of feedback instances (Feedback). The values in parentheses correspond to a single standard deviation over 10 trials.
Grid Transportation-Easy Grid Transportation-Hard
Success Steps Feedback Success Steps Feedback
2$\times$2 HMAS-2 100% 9.9(2.74) 3.3(2.05) 80% 7.0(5.0) 6.0(9.74)
LLaMAC 100% 7.0(1.79) 2.0(1.26) 100% 4.7(1.35) 3.6(2.80)
2$\times$4 HMAS-2 80% 15.5(6.09) 12.3(5.83) 20% 17.0(9.0) 24.0(20.0)
LLaMAC 100% 7.6(1.36) 4.3(1.42) 90% 7.44(2.95) 10.56(7.54)
4$\times$8 HMAS-2 60% 30.6(9.70) 26.1(13.59) 0% - -
LLaMAC 100% 12.9(2.70) 10.7(3.35) 90% 8.44(1.57) 12.11(2.51)
As shown in Figure 4, it is evident that within a limited number of steps, LLaMAC demonstrates the ability to explore and learn through continuous interaction with the environment. The final performance of all methods is presented in Figure 5. The TripletCritic approach within LLaMAC exhibits a similar structure to the Multi-agent Debate method, and compared to other approaches, these two methods display relatively stable performance. However, debate-based methods often suffer from excessive or insufficient exploration, resulting in a tendency to converge to local optima. On the other hand, approaches that emphasize exploration and exploitation struggle to maintain stable performance. The former exhibits significant oscillations due to excessive exploration, while the latter prematurely converges to local optima after only a few simple exploratory steps, aligning with the expected characteristics of these methods. Distributed approach incurs the highest access cost , as each agent is required to independently access the LLM. Nevertheless, the lack of collaboration among the agents still hinders the capture of true relationships.
@@ 4.2.1 Experimental Settings
4.2.2 Results
Figure 7: Token usage of LLaMAC and HMAS-2 in the Grid Transportation scenarios.
We conduct a comparative analysis between our method and the state-of-the-art solution, HMAS-2 Chen et al. (2023b). For each scenario, we conduct tests on grid configurations of $2\times 2$, $2\times 4$, and $4\times 8$, respectively. Table 2 presents a comprehensive performance comparison between the two methods, clearly demonstrating the overall superiority of our approach. In complex scenarios involving long-term iterative decision-making, LLaMAC exhibits a significantly higher success rate compared to HMAS-2. Furthermore, LLaMAC consistently achieves task completion in fewer interaction steps, highlighting the performance advantages of its employed strategies. Additionally, as shown in Figure 7 the TripletCritic facilitates the generation of superior initial suggestions, thereby reducing the need for feedback iterations and greatly enhancing token utilization efficiency.
4.2.3 Case Study
During the experimental process, we observe that LLaMAC effectively enhances the capabilities of LLM in long-term planning and execution, spatial reasoning, and learning from interactions or errors. For example, spatial reasoning poses a significant challenge for LLMs, as they are more prone to hallucinations when determining whether an object is closer to the target. This issue becomes more pronounced in the Hard scenario. As shown in Figure 8, in the HMAS-2 method, agents often move objects to positions far from the target and may repeatedly move them between two particular locations. In contrast, in LLaMAC, such occurrences are often corrected during the external feedback phase. The actor only needs to focus on its own task, and when it receives suggestions from the critic, the difficulty of determining the effectiveness of individual agent tasks is significantly reduced compared to joint policies. This makes spatial reasoning errors more easily detected, reflected and corrected.
@@ 5 Conclusion
In this study, we present a novel framework called LLaMAC to enhance the collaborative performance of large-scale multi-agent systems based on Large Language Models. Building upon the commonsense reasoning capabilities exhibited by LLMs, we effectively augment the planning and coordination abilities among agents through stable reasoning mechanisms and comprehensive feedback mechanisms, facilitating continuous interaction between agents and the environment. LLaMAC demonstrates remarkable performance in coordinated scenarios involving a large number of agents. Notably, it exhibits exceptional capabilities in long-term planning, mathematical reasoning and optimization problems, spatial reasoning, and learning from mistakes. Additionally, LLaMAC reduces the access costs associated with large-scale multi-agent collaboration. We believe that with further enhancements in LLMs and the emergence of more collaboration frameworks, the field of multi-agent collaboration will experience new opportunities for advancement.
Appendix A Implementation Details
A.1 Pseudo-Code
Algorithm 1 Execution Procedure for LLaMAC
Hyperparameters: Length of episode $T$, number of agents $N$, trajectory length of Memory $L$, maximum number of internal and external feedback iterations $IF,EF$
Initialize: Memory $\mathcal{M}$, Environmental initial state $s_{0}$ and observation $\{o_{0}^{i}\}_{i\in\mathcal{I}}$, timestep $t=0$
1:  while $t\leq T$ do
2:     TripletCritic receives the memory information $m_{t}$ and the current state $s_{t}$
3:     Generate suggestion for all actors $su=\{su_{t}^{i}\}_{i\in\mathcal{I}}$ through Internal Feedback (Algorithm 2)
4:     Genearate the joint action $\mathbf{a}_{t}=\{a_{t}^{1},a_{t}^{2},\dots,a_{t}^{n}\}$ through External Feedback (Algorithm 3)
5:     Execution the joint action $\mathbf{a}_{t}$, obtain reward $r^{i}_{t}$ and environmental state $s_{t+1}$
6:     Collect trajectories $\tau^{i}$, push transitions $\{(s_{t},a^{i}_{t},r^{i}_{t},s_{t+1}\}$ into $\mathcal{M}$
7:  end while
Algorithm 2 Internal Feedback
[…]
1:  while $f_{i}\leq IF$ do
2:     for critic $j=1$ to $2$ do
3:        Generate actions $\mathbf{a}_{j}=\{a_{t}^{i}\}_{i\in\mathcal{I}}$ corresponding to preference $\mathbf{a}_{j}\sim\mathbf{LLM}_{critic_{j}}(m_{t},s_{t},F_{if})$
4:     end for
5:     Assessor makes Veracity Scrutiny, get results $vs$
6:     if $vs$ is True then
7:        Assessor makes Belief Correction, generate final action suggestion for all actors $su=\{su_{t}^{i}\}_{i\in\mathcal{I}}$, where $su\sim\mathbf{LLM}_{assessor}(m_{t},s_{t},\mathbf{a}_{1},\mathbf{a}_{2})$
8:        break
9:     else
10:        Generate feedback information $F_{if}$
11:     end if
12:     $f_{i}=f_{i}+1$
13:  end while
Algorithm 3 External Feedback
Input: Maximum number of Enternal feedback iterations $IF$, current iteration number $f_{e}=0$, feedback information $F_{ef}=[]$, suggestion from TripletCritic $su$, observation $\{o_{t}^{i}\}_{i\in\mathcal{I}}$, state $s_{t}$
1:  while $f_{e}\leq EF$ do
2:     for agent $i=1$ to $N$ do
3:        Actor $i$ makes Plan Confirmation
4:        if Execute then
5:           $a_{t}^{i}=su_{t}^{i}$
6:        else
7:           Generate actor feedback Information $F^{i}_{ef}\sim\mathbf{LLM}_{actor_{i}}(o^{i}_{t},su^{i}_{t})$
8:           $F_{ef}=F_{ef}+F^{i}_{ef}$
9:        end if
10:     end for
11:     if $F_{ef}$ is not [][][ ] then
12:        Assessor regenerates action suggestions $su\sim\mathbf{LLM}_{assessor}(m_{t},s_{t},F_{ef})$
13:     else
14:        break
15:     end if
16:     $f_{e}=f_{e}+1$
17:  end while
A.2 Prompt Example in llamac
As shown in Section A.1, the entire process of LLaMAC’s iterative decision-making is facilitated by internal and external feedback mechanisms, enabling seamless collaboration among its modules to accomplish decision tasks in large-scale intelligent agent systems. As shown in Figure A.1 and Figure A.2, we demonstrate the core content of the prompt used by LLaMAC to access LLM during the decision-making process.
Figure A.1: Prompt for the large language model in the internal feedback mechanism.
Figure A.2: Prompt for the large language model in the external feedback mechanism. Only actors that are determined by the plan confirmation to be reused will execute this access.
Appendix B Environment details
B.1 System Resource Allocation
The system resource allocation environment can be regarded as an optimization problem or a single-step decision problem, where the available actions of all agents are fixed at each decision-making instance. The memory stores the observation history of the agents in the form of a dictionary: [{action:[], system_reward:[]}, …, {action:[], system_reward:[]}]. Additionally, we require the decision-makers to simultaneously output thoughts and actions to enhance the reasoning capability of the language model.
B.2 Grid Transportation
Grid transportation tasks are inherently more complex and demand higher decision-making capabilities. They involve language models assuming different roles to collaborate through continuous dialogue and interaction, generating long-term action trajectories, and ultimately achieving the final objectives.
In this environment, the Veracity Scrutiny within the Internal Feedback involves policy checks of the joint strategy and is set to evaluate (1) whether the output grammar conforms to the specified format and (2) whether the joint actions result in conflicts. The Plan Confirmation within the External Feedback involves policy checks specific to each agent and is set to evaluate (1) the availability of actions and (2) whether the suggestions result in a shorter Manhattan distance between objects and targets. Taking the Hard scenario as an example, the variables utilized in the decision-making process of the intelligent agents are depicted in Figure B.1.
Figure B.1: Text description of states, observations, and actions in the grid transport environment.